strutture:lnf:da:plasmonx:lpa:spectrometer:v1commissioning:jul2012
July 10th 2012
| Run | time | run conditions | comments |
| 760 | 12:30 | pedestals | ch 0,1,2 with PMTs |
| 761 | 12:40 | LED | HV on |
| 762 | 13:00 | LED | HV on, change timing |
| 763 | 18:35 | electrons but not to chamber | HV 900V |
| 764 | 19:15 | electrons but not to chamber | HV 500V |
| 765 | 19:45 | electrons but not to chamber | HV 500V, ch0 light tight, flanges darkened |
| 766 | 19:57 | electrons but not to chamber | HV 500V, ch0 light tight |
| 767 | 19:57 | electrons but not to chamber | HV 500V, ch0 light tight, dt=150 |
| 768 | 19:57 | electrons but not to chamber | HV 500V, ch0 light tight, dt=250 |
| 769 | 20:21 | electrons but not to chamber | HV 500V, ch0 light tight, dt=50, magnet on |
| 771-781 | 22:31 | problems after powercut | trigger timing changed |
| 782 | 0:21 | electrons in chamber | HV 500V, all light tight, dt=50, B=0 |
| 783 | 0:21 | electrons in chamber | HV 500V, all light tight, dt=50, B=0.45 |
July 11th 2012
HW changes:
1) the second stage of the dual timer has been replaced by a Stanford for reliable and reproducible setting. 2) all five channels have been equipped with PMT+Electronics 3) the chamber of the spectrometer has been connected with the interaction region (i.e. the glass window removed) 4) a lead collimator has been inserted between the gas jet and the spectrometer
Plan of the day:
A) scan in integration time to find the best timing [with small magnet in] B) scan in HV to equalize gains [with small magnet in] C) spectrum with B=0 and B>0 (with small magnet out, having subtracted the common background)
| Run | time | run conditions | comments |
| 789 | 15:47 | No Beam HV on ch 0-4 500 V | pedestals |
| 790 | 16:22 | No Beam HV on ch 0-4 900 V | pedestals |
| 793 | 16:22 | electrons | HV=500V. EM Noise is Back! |
| 794-797 | 17-18 | Trigger Test | the trigger doesn't works |
| 798 | 19:38 | T (signal splitting) removed from trigger | the trigger works |
| 799 | 19:44 | HV = 0V, laser+plasma | Electronic noise |
| 800 | 20:00 | HV=0, laser+plasma, reinforce grounding | |
| 801 | 20:04 | try HV=500 but impossible to take data | |
| 804 | 22:50 | HV=700 PMT 0, HV=0 PMT 1-4 | saturation on PMT 0 |
| 805 | 22:52 | HV=600 PMT 0 1 2, HV=0 PMT 3 4 | now verify that noise on PMT0 increases with num of connected cards |
| 806 | 23:26 | HV=600 PMT 0 1 2, HV=0 PMT 3 4 | only card 1 is connected, with the attenuator |
| 807 | 23:36 | HV=600 PMT 0 1 2, HV=0 PMT 3 4 | B=0.45T |
| 808 | 23:47 | HV=600 PMT 0 1 2, HV=0 PMT 3 4 | B=0.45T, card 1 2 3 are connected with attenuators |
| 809-814 | 23:59 | HV=600 PMT 0 1 2, HV=0 PMT 3 4 | B=0, scan in time reveals that t=0 is the best –>anticipate trigger |
| 816-821 | 24:10 | HV=600 PMT 0 1 2, HV=0 PMT 3 4 | B=0, scan in time: best timing dt = 100 ns |
| 822 | 24:27 | HV=550 V PMT 0, HV=600 PMT 1 2, HV=0 PMT 3 4 | B=0 |
| 823 | 24:30 | HV=550 V PMT 0, HV=600 PMT 1 2, HV=0 PMT 3 4 | B=0 |
| 824 | 24:37 | HV=500 V PMT 0, HV=600 PMT 1 2, HV=0 PMT 3 4 | B=0 |
| 825 | 24:37 | HV=OFF | B=0 |
| 826 | 24:41 | HV=900 PMT 0 1 2 | B=0 |
| 827 | 24:45 | HV=900 PMT 0 1 2, dt=0 | B=0 |
July 12th 2012
Investigated noise problem. Disappears by using screened cable for DY12
| Run | time | run conditions | comments |
| 829 | 13:29 | hv 800 ch 0,1,2 | B=0 |
| 830 | 13:32 | hv 900 ch 0,1,2 | B=0 |
| 831 | 13:35 | hv 800 ch 0,1,2 | B=0 magnete flame in |
| 832 | 14:10 | hv 500 ch 0,1,2 | B=0 magnete flame in remove attenuators |
| 833 | 14:15 | hv 600 ch 0,1,2 | B=0 magnete flame in remove attenuators |
| 834-852 | 15:33 | hv 600 ch 0-4 | trigger problems |
| 853-858 | 16:40 | hv 600 ch 0-4 | time scan |
| 859-870 | 16:45 | hv 500 ch 0-4 | time scan |
| 871-872 | 17:30 | HV scan to equalize | establish that HV={580,550,500,480,480}. Found that last two cables are swapped |
| 873 | B>0 | turn on magnet. Hv default. | Elettroni deviati!!! |
| 874 | B>0 | magnet ON. HV default | costanti di calibrazione Offline C=(1.2,1,1.5,1.1,1.4) |
| 875 | B>0 | HV def | Collimatore: 2 blocchetti 5 cm Pb dopo la finestra con 2 mm separazione: 40 mrad * 100 cm –> dx=4 cm |
| 876 | B=0 | HV def | Collimatore; B ramping down? |
| 877 | B=0 | HV def | Collimatore |
| 878 | B=0 | HV def | Collimatore |
| 879 | B=0 | HV def | Collimatore chiuso; scambiati canali 3 e 4 che erano invertiti; C=(1.2,1,1.5,1.4,1.1) |
| 882 | B=0 | HV def | Collimatore chiuso |
| 883 | B=0 | HV def | Collimatore chiuso; Flame Magnet IN; Use for Ped File |
| 884 | B=0 | HV def | Collimatore aperto 2mm; Flame Magnet OUT; usb problem at the end |
| 885 | B=0 | HV def | Collimatore aperto 2mm; Flame Magnet OUT |
| 886 | B=0 | HV def | Collimatore aperto 1cm; Flame Magnet OUT |
| 887 | B=0.5 T | HV def | Collimatore aperto 1cm; Flame Magnet OUT |
July 13th 2012
Debug of spectrometer trigger. Problems with upper part of logic unit, replaced with lower part. Dual timer removed, veto implemented with Stanford D signal set to 500 mus. Delay and width of trigger button revised, set to 500 mus. Missing alignment with photocatode.
July 18th 2012
Pedestal studies
| 896 | 12:21 | HV off | Only maroc 0 board+cable; no laser | |
| 897 | 12:26 | HV off | Only maroc 0 board+cable; no laser | RMS 12 cts |
| 899 | 12:26 | HV off | Only maroc 0 board no cable; no laser; improved trigger cable | RMS 1.5 cts |
Giulietto Felici works on electronics to improve grounding and box.
July 25th 2012
test electronics with new shielding
| 900 | 10:37 | pedestals | 5cards but no cables; RMS 1-2 ADC |
| 901 | 10:40 | pedestals | 5cards cable #0 in; HV off; RMS ~6 for first 43 channels, 10-18 for the others |
| 902 | 10:44 | pedestals | 5cards cable #0 in; HV=900V; RMS ~6 for first 43 channels, 10-18 for the others |
| 903 | 10:52 | LED run | 5cards cable #0 in; HV=900V |
| 904 | 14:26 | pedestals | 5cards all cables in; HV=0V |
| 905 | 14:29 | pedestals | 5cards all cables in; HV=default |
| 905 | 14:29 | pedestals | 5cards all cables in; HV=default – high stat |
strutture/lnf/da/plasmonx/lpa/spectrometer/v1commissioning/jul2012.txt · Last modified: 2012/07/25 12:30 by faccini@infn.it
